what is the substitution method of y=2x-1 and y=-3x+4
Answer by Alan3354(69443)

what is the substitution method of y=2x-1 and y=-3x+4
------------------
The substitution method can be used for any set of equations.
I'll proceed assuming you meant to ask how the substitution method can be used to solve for x and y in the 2 equations you entered.
------------------------
y=2x-1 and y=-3x+4
Since both expressions on the right sides of the eqns = y, they equal each other.
2x-1 = -3x+4
This is substitution, the right sides were substituted for y.
Add 3x
5x-1 = 4
Add 1
5x = 4
x = 1 xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx
Now substitute 1 for x into either eqn to find y.
y = 2*1-1
y = 1 yyyyyyyyyyyyyyyyyyyyyyyyyyyy
we're done.